OpenStudy (anonymous):

Solve 6/x-4 = 4/x and determine if the solution is extraneous or not. Medal given! please help!

OpenStudy (anonymous):

@zepdrix @campbell_st

zepdrix (zepdrix):

\[\Large\bf\sf \frac{6}{x-4}\quad=\quad \frac{4}{x}\]Is this the problem?

OpenStudy (anonymous):

yes

zepdrix (zepdrix):

So ummm how should we approach this... Remember how to cross multiply?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

would it be... 6x = (x-4)(4)

zepdrix (zepdrix):

mmm ok good

zepdrix (zepdrix):

So we need to distribute the 4 to each term in the brackets, yes?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

(4)(x-4) = 4x-16

OpenStudy (anonymous):

the choices are:

zepdrix (zepdrix):

woops what happened in the middle there... 6x-4x= 16 2x = 16, right?

OpenStudy (anonymous):

x= -8, extraneous x= -8, non-extraneous x= 8, extraneous x= 8, non- extraneous

OpenStudy (anonymous):

woops, its 8

OpenStudy (anonymous):

typo

zepdrix (zepdrix):

Extraneous means BAD! If we plug 8 back into the original problem, does it cause a problem? Does it cause us to divide by zero in any spot?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

so non-extraneous

zepdrix (zepdrix):

Yay good job!

OpenStudy (anonymous):

ugh that was easy.. I could've done that one myself... could you help me with a harder one like that

zepdrix (zepdrix):

Oooo we made a little boo boo hold on!

OpenStudy (anonymous):

oh no where

zepdrix (zepdrix):

\[\Large\bf\sf 6x\quad=\quad 4x-16\]Subtracting 4x to each side should have given you,\[\Large\bf\sf 6x-4x\quad=\quad -16\]\[\Large\bf\sf 2x\quad=\quad -16\]

OpenStudy (anonymous):

ohh

OpenStudy (anonymous):

so -8

OpenStudy (anonymous):

non-extraneous

zepdrix (zepdrix):

mm k ya i think we've got it now! c:

OpenStudy (anonymous):

ok here's my last one

OpenStudy (anonymous):

can u help me quickly, i need to do something in a few minutes before i leave the house :/

zepdrix (zepdrix):

\[\Large\bf\sf \frac{6}{x-6}\quad=\quad \frac{x}{x-6}-\frac{6}{2}\]There's a few different ways to do this one. See how we have a couple fractions with the same denominator? Let's combine those by getting them on the same side. So we'll subtract x/(x-6) from each side.\[\Large\bf\sf \frac{6}{x-6}-\frac{x}{x-6}\quad=\quad -\frac{6}{2}\] \[\Large\bf\sf \frac{6-x}{x-6}\quad=\quad -\frac{6}{2}\]

zepdrix (zepdrix):

Then from there you can cross multiply again I guess.

OpenStudy (anonymous):

wait! stay here

OpenStudy (anonymous):

(2)(6-x)= (-6)(x-6)

OpenStudy (anonymous):

12-2x = -6x+36

zepdrix (zepdrix):

k looks good so far!

OpenStudy (anonymous):

-2x+6x= -12+36

OpenStudy (anonymous):

-4x= 24 x= -6

OpenStudy (anonymous):

do you think its extraneous or non?

zepdrix (zepdrix):

\[\Large\bf\sf -2x+6x\quad\ne\quad -4x\]

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

4x?

zepdrix (zepdrix):

ya

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

4x= 24 x= 6

OpenStudy (anonymous):

so is it extraneous or non-extraneous

zepdrix (zepdrix):

Hmm does it cause any problems when we try to plug it back into the original?\[\Large\bf\sf \frac{6}{\color{red}{6}-6}\quad=\quad \frac{\color{red}{6}}{\color{red}{6}-6}-\frac{6}{2}\]

OpenStudy (anonymous):

yea

zepdrix (zepdrix):

Hmm ya, looks like it's giving us zeros in our denominators. Bad! Extraneous! :O

OpenStudy (anonymous):

thanks so much! i gtg now! :)

zepdrix (zepdrix):

bb \c: